OpenStudy (jadenwalker76):

Choose the equation below that represents the line passing through the point (−3, −1) with a slope of 4. y − 1 = 4(x − 3) y + 1 = 4(x + 3) y − 3 = 4(x + 1) y + 3 = 4(x − 1)

OpenStudy (imstuck):

Are you familiar with the slope intercept form of a line? y = mx + b?

OpenStudy (imstuck):

What you do in this type of a situation is to plug in what they give you and solve for b, the y intercept, and the go back into the general form, plug the slope back in that they give you to be 4 and the b that you solved for, and there you have it!

OpenStudy (imstuck):

In the point they gave you, y is -1 and x is -3, and you have a slope of 4. Plug these in like this: y = mx + b -1 = 4(-3) + b and solve for b.

OpenStudy (imstuck):

When you get a solution for b, plug it back in like this: y = 4x + ____

OpenStudy (imstuck):

It could also be a negative solution, and in that case, you'll have y = 4x - ____

OpenStudy (imstuck):

Oh I see you need it in point slope. No worries. That's even easier!

OpenStudy (imstuck):

This is point-slope form for a general line: \[(y-y _{1})=m(x-x _{1})\]

OpenStudy (imstuck):

Again you have a y value from the point of -1 and an x value of -3, with a slope of 4. Just plug in the numbers.

OpenStudy (imstuck):

Keep in mind the rules of subtracting negative numbers from each other.